OpenStudy (anonymous):

consider the equation x square + 10x = 15 a) what # must be added to x square + 10 to complete the square? b) solve the equation x square + 10 =15 by adding that # on both sides n using the method shown in example one c) solve the quation x square + 10x = 15 using the quadratic formula

OpenStudy (maya):

add 25 on left hand side

OpenStudy (maya):

so you get (x+5)^2=15

OpenStudy (anonymous):

\[x^2+10x=15\] to complete the square think half of ten is five then write \[(x+5)^2=15+\] and \[5^2 = 25\] to get \[(x+5)^2=15+25=40\]

OpenStudy (anonymous):

so now you have \[(x+5)^2=40\] and \[x+5=\pm \sqrt{40}\]so \[x=-5\pm\sqrt{40}\]

OpenStudy (anonymous):

if you have to write in simplest radical form \[\sqrt{40}=\sqrt{4\times 10}=\sqrt{4}\times \sqrt{10}=2\sqrt{10}\]

OpenStudy (anonymous):

so 'final answer" is \[-5\pm2\sqrt{10}\]